OpenStudy (anonymous):

2 balls are chosen randomly from an urn containing 8 white, 4 black, 2 orange ball. Suppose you win 2 dollars for each black ball and lose 1 dollar for each white ball selected. Let X be the net winning , Find a) S_x my answer is for a) (-2,1,2)

OpenStudy (anonymous):

b) asked for pmf P_x (8/14)^2 x=-2 (4/14)*(8/14) x=1 (4/14)^2 x=2

OpenStudy (anonymous):

@zarkon

OpenStudy (anonymous):

@myininaya

OpenStudy (mathlegend):

What type of math is this? Algebra 1?

OpenStudy (mathlegend):

Honestly, I have no clue.

OpenStudy (openstudyuser):

lol

OpenStudy (anonymous):

probability

OpenStudy (zarkon):

sampling with or without replacement?

OpenStudy (anonymous):

didn't say anything about replacement so I think without

OpenStudy (zarkon):

if it is without...then your pmf is not correct

OpenStudy (anonymous):

what is it mean by "replacement"?

OpenStudy (zarkon):

if it is replacement then you pick a ball...look at it ..put it back then randomly sample again with all 14 balls

OpenStudy (anonymous):

I think ,from the wording; we are pulling two balls out at the same time

OpenStudy (zarkon):

that is what I am thinking

OpenStudy (zarkon):

so you need to redo your pmf

OpenStudy (zarkon):

if you get black balls then you get $2 each correct

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

so one possibility is that you win $4

OpenStudy (anonymous):

white ball, I lost $1

OpenStudy (zarkon):

BB=$4 BO=$2 OB=$2 BW=$1 WB=$1 ...

OpenStudy (zarkon):

so X can take the values 4,2,1,0,-1,-2

OpenStudy (zarkon):

agree?

OpenStudy (anonymous):

oh, I see, I omitted orange which I shouldn't have

OpenStudy (anonymous):

so how would I write PMF?

OpenStudy (zarkon):

what i s the probability of 2 blacks? ie $4

OpenStudy (anonymous):

(8/14)*(7/13)

OpenStudy (zarkon):

ok what about winning $2

OpenStudy (anonymous):

(4/14)(2/13) or (2/14)(4/13) ?

OpenStudy (zarkon):

winning $2 is BO or OB

OpenStudy (zarkon):

there are 8 blacks...i see no 8's in your solution

OpenStudy (zarkon):

lol.....there are 8 W

OpenStudy (zarkon):

4 black...so prob of BB is \[\frac{4}{14}\cdot\frac{3}{13}\]

OpenStudy (zarkon):

not what you had above

OpenStudy (zarkon):

so \[P(X=4)=\frac{4}{14}\cdot\frac{3}{13}\]

OpenStudy (anonymous):

right , so for $2 (4/14)(2/13) or (2/14)(4/13) (4/14)(2/13)+(2/14)(4/13)=

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

the sum of the two

OpenStudy (anonymous):

I understand now; I really appreciate your help

OpenStudy (zarkon):

no problem